OpenStudy (anonymous):

Multiply: 2 the square root of 11 • 8 the square root of 22

OpenStudy (anonymous):

2 sqrt 11 • 8 sqrt 22 = 2(8) • sqrt (11 • 22) = 16 sqrt 242 = 16 sqrt (121 • 2) = 16 sqrt 121 sqrt 2 = 16(11) sqrt 2 = 176 sqrt 2

OpenStudy (akashdeepdeb):

\[2 \sqrt 11 • 8 \sqrt 22 = 2(8) • \sqrt (11 • 22) = 16 \sqrt 242 = 16 \sqrt (121 • 2) = 16 \sqrt 121 \sqrt 2 \] \[ = 16(11) \sqrt 2 = 176 \sqrt 2\]